# Synthetic Biology Meets Neuromorphic Computing: Towards a Bio-Inspired Olfactory Perception System

### v1

@@ 4 Conclusions & Outlook
While process variations and mismatch in weak-inversion circuits may introduce variability, the system integrates multiple biasing mechanisms to compensate for these effects, which will be further validated during hardware testing. The bigger source of drift and variability might stem from the biological receptors, for example by leakage of ions from the synthetic cell vesicles. This will make it necessary to adapt electrical methods for compensation, e.g. DC drift offset compensation, or more systematic approaches like homeostatic learning mechanisms.
@@
We plan to extend the current work with hardware characterization, comprehensive dataset collection, and system scalability. Specifically, efforts will be directed toward expanding the platform to support multiple sensor pads, each featuring distinct odor receptor chambers, thereby enhancing the system’s capability to distinguish between a wider range of chemical compounds. In addition, spiking neural network classifiers will be used to achieve real-time classification of complex odor mixtures, paving the way for practical applications in autonomous chemical analysis.
The results of our model pipeline give us strong guardrails for the development of our platform, informing the co-design of Orco-OR complexes, vesicle properties, amplification stages, and decoding hardware. As seen in Fig. 6, the classification of the odor is confounded by the noise and the number / types of available OR. In this regard, we interpret the SNN classification results as a baseline.
@@
In addition, as part of our co-design, we are developing network models which are able to extract relevant information delivered from the sensors, such that high classification accuracy with minimal number of ORs can be achieved. A crucial data feature may be the temporal structure of the stimulus. The hardware simulations shown here are based on the direct injection of the sensor current into the first layer of neurons, effectively resulting in rate coding with a bin width of $\mathcal{O}(200)\,$ms (Fig. 8). By adapting circuit components, we are able to change the time constants to fully exploit the temporal structure of the stimulus. Alternatively, coding schemes such as precise spike timing have been observed in our model species [96], and may prove valuable for our application.
@@
The setup described here is limited to supervised training using single odorants. Future work will investigate odor mixtures as they appear in real-world applications. Due to OR co-activation, it is unclear which response to expect from the synthetic cells, and careful characterization of the odor mixtures will be needed. Finally, we aim to investigate unsupervised and few-shot learning for real-world use of our sensing platform.
@@ 5.3 Electrode model
Multielectrode Arrays (MEAs) consisting of 60 titanium nitride electrodes of 30 $\mu$m diameter (Model 60tMEA200/30iR-ITO-gr, Multichannel Systems, Germany) were characterized using impedance spectroscopy. Impedance spectra in the frequency range 10 kHz to 1 MHz were obtained in phosphate buffer solution of physiological salinity at room temperature. A three-circuit model consisting of a resistor in series with capacitance ($C_{dl}$) and a resistor ($R_{t}$) in parallel was used to fit the spectra with $\chi^{2}\approx 0.2$. The resistance of the gap $R_{gap}$ was calculated from $R_{gap}=\rho_{s}/(5\pi d_{gap}$), with approximate values for the solution of conductivity $\rho_{s}=1$ $\Omega$m and $d_{gap}=1$ nm [93, 99]. The final values for the electrode model are $R_{\text{gap}}=6.37\times 10^{7}\,\Omega$, $C_{\text{dl}}=8.99\times 10^{-11}\,\text{F}$, and $R_{\text{t}}=7.92\times 10^{8}\,\Omega$.
@@ 5.4 Odor space analysis
We describe our method for selecting the most viable ORs for a given set of odorants in Algorithm 1. The method requires a response data matrix with entries (odor $\times$ OR), which we have retrieved from the DoOR 2.0.1 database [70, 71]; specifically, we use the provided door_response_matrix.csv, which contains responses of 78 sensory unit types to 692 odorants, normalized to $[0,1]$ across all responding units. We first exclude all unit types not clearly linked to a single OR type, as described in [71], which leaves a total of 52 ORs to be analyzed. As only a subset of all unit type/odor combinations have been measured (about 14 %), we replace missing entries (NA) with zero response. This may introduce the issue that some odors are separated only by the absence of any response; in practice, this would make it impossible to distinguish such an odor from noise currents when no ligand is present. We mitigate this issue by including a ‘no ligand/none’ odor in the analysis, see below.
@@
Our algorithm for selecting odor/OR sets is easily implemented and performs the reduction quickly. Note however that the underlying dataset corresponds to response measurements of olfactory receptor neurons corrected for the baseline firing rate, and doesn’t cover cases where no ligand is present at all, as described above. Therefore, if the final selection given by our algorithm includes odors for which none of the selected ORs respond at all, this odor will be experimentally indistinguishable from no ligand binding. To cover this case, we have included a ‘no ligand/none’ odor, i.e. a row in $D_{\text{all}}$ which has zero response for all OR.
@@ 5.6 Design and Simulation of the Neuromorphic Hardware Platform
The system integrates an analog front-end with an event-driven, asynchronous spiking neural network and a microcontroller. It consists of: (i) a bioelectronic interface coupling lipid vesicles with synthetic olfactory receptors and microelectrode arrays, (ii) a low-noise analog amplifier for signal conditioning, (iii) a spiking neural network for energy-efficient sensor response processing, and (iv) a RISC-V microcontroller for spike decoding and synaptic weight configuration. Figure 9 illustrates the layout of the full chip. The low-noise amplifier (LNA) exploits a folded cascode OTA (FC-OTA) with an active cascode and a flipped current follower (FCF) to improve signal integrity, transconductance efficiency, and noise performance. The FC-OTA topology enhances gain while maintaining a wide output swing, the active cascode increases output impedance for better gain stability, and the FCF reduces impedance at the folded node, ensuring efficient current transfer. A detailed overview of these techniques is provided, with further details in the work by R. Sanjay et al. [101]. Somatic and synaptic circuits are designed with weak-inversion transistors and follow the principles of well-known neuromorphic designs [102, 103] that have been demonstrated to scale to advanced technologies nodes [104, 105].

### v2

@@ 4 Discussion & Outlook
In addition, while process variations and mismatch in weak-inversion circuits introduce variability in neuromorphic systems [111] our system integrates multiple biasing mechanisms to compensate for these effects, which will be further validated during hardware testing. However, it is worth noting that bio-inspired computational strategies often exploit the variability in all the network parameters as a computational resource for example in learning applications [112], or in dynamical computation using attractor networks [113] and decision perceptual making [114] tasks; we plan to test these computational primitives in the lab with real chemosensory data. In our system, we expect the bigger source of drift and variability to stem from the biological receptors, for example by leakage of ions from the synthetic cell vesicles which has a typical timescale of days  [115]. This could make it necessary to adapt electrical methods for compensation, e.g. DC drift offset compensation, or more systematic approaches like homeostatic learning mechanisms [116].
@@
While the actual hardware implementation of hybrid synethtic sensory neurons is underway, we plan to extend the current work with hardware characterization, comprehensive dataset collection, and demonstration of system scalability. Specifically, efforts will be directed toward expanding the platform to support multiple sensor pads, each featuring distinct odor receptor chambers, thereby enhancing the system’s capability to distinguish between a wider range of chemical compounds, and our mixed-signal spiking neural network hardware classifiers will be used to achieve real-time classification of complex odor mixtures.
One important stipulation of our simulation pipeline are the underlying odorant concentrations levels. To model odor responses, we have used the DoOR dataset, which contains recordings with high concentrations of volatile compounds, orders of magnitude higher than naturally occuring concentrations. While this increases the signal-to-noise ratio in spike recordings, it also reduces the selectivity of individual ORs [27, 30]. I.e., there is a trade-off between sensitivity and selectivity in neuron responses, and we expect this to be present also in our hybrid synthetic sensory neurons. In a real-world application of our sensing platform with natural concentrations, this trade-off is likely to manifest as a lower signal-to-noise ratio, but tighter matching of individual ORs to odors. It is therefore of high interest to repeat the odor space analysis under natural and artifically high concentrations using our sensing system. This will not only help us fine-tune the performance of our platform, but may also provide insights about the sensitivity/selectivity curve of sensory neurons in vivo.
@@
In addition, as part of our co-design, we are developing network models which are able to extract relevant information delivered from the sensors, such that high classification accuracy with minimal number of ORs can be achieved. A crucial data feature may be the temporal structure of the stimulus [117]. The hardware simulations shown here are based on the direct injection of the sensor current into the first layer of neurons, effectively resulting in rate coding with a bin width of $\mathcal{O}(200)\,$ms (Fig. 8). By adapting the bias currents that control the parameters of the synapses and neurons circuits, such as time constants, leak currents, etc., we aim to fully exploit the temporal structure of the stimulus. In parallel, coding schemes such as precise spike timing have been observed in our model species [118], and may prove valuable for our application.
@@
The setup described here is limited to supervised training using single (non-mixed) odor presentations. Future work will investigate odor mixtures as they appear in real-world applications. In this regard, it may prove valuable to model ligand-OR interactions as fuzzy sensors, as proposed in [119]. Due to OR co-activation, it is unclear which response to expect from the synthetic cells under odor mixtures, and careful characterization will be needed. Finally, we aim to investigate unsupervised and few-shot learning for real-world use of our sensing platform, following recent progress in understanding continual learning in the fruit fly olfactory system [69].
Moving beyond insect olfaction, recent advances have also demonstrated successful integration of human olfactory receptors (hORs) with organic neuromorphic synaptic devices, achieving highly accurate, molecule-specific odor discrimination [120]. Unlike our approach, which employs synthetic insect olfactory receptor complexes coupled to analog CMOS spiking networks, Song et al. [120] used part of the human olfactory system interfaced with organic electronics, showing precise discrimination of structurally similar odorants. Although their work highlights the feasibility of exploiting human olfactory specificity for targeted sensing applications, our framework emphasizes modularity and adaptability through synthetic biology and neuromorphic codesign, aiming to address broader olfactory detection challenges. Finally, Wu et al. [121] proposed a bionic olfactory neuron by combining organic field-effect transistor arrays with in-sensor reservoir computing to achieve classification for complex odor mixtures, isomers, and homologs, however, our methodology is complementary as we uniquely integrates synthetic biology-based receptors with neuromorphic circuits, potentially enabling direct biological specificity and real-time adaptability.
5 Conclusion
This work presents a novel co-design methodology of a neuromorphic chemosensing system that aims at integrating biological olfactory receptors with analog neuromorphic circuits, mimicking natural olfactory processing. We demonstrate, using computer simulations, the use deep-learning tools such as AlphaFold to select suitable odorant receptors (ORs) and we generate a synthetic dataset in which each Orco-OR complex is modelled as an independent two-state Markov process. The resulting open/close transitions yield continuous ion-channel currents that our electrode model converts into analogue voltage traces (Fig. 5). These noisy, microvolt-level signals feed directly into an NGSpice simulation of the neuromorphic front-end (Fig. 8), where sub-threshold analogue neurons transform them into spikes. A spiking neural network, supervisely trained on the same odors, then classifies the spike patterns with high accuracy (Fig. 6). Importantly, by combining biological odor receptors with neuromorphic hardware, we aim to achieve real-time, energy-efficient chemosensory detection. The use of weak-inversion transistors and event-driven spiking neural networks ensures on-demand processing, minimizing power consumption while maintaining high sensitivity to chemical stimuli.
Our approach is unique in that it combines the co-design of the physical realization of sensory neurons, processing, and learning mechanisms, all in one device. To reach this goal, we build on recent advances in the use of machine learning methods in biology to enable the co-design with established electrical engineering methods.
The results of our model pipeline give us strong guardrails for the development of our platform, informing the co-design of Orco-OR complexes, vesicle properties, amplification stages, and decoding hardware. As seen in Fig. 6, the classification of odors is confounded by the noise and the number / types of available OR. In this regard, we interpret the SNN classification results as a baseline.
@@ 6.3 Electrode model
Multielectrode Arrays (MEAs) consisting of 60 titanium nitride electrodes of 30 $\mu$m diameter (Model 60tMEA200/30iR-ITO-gr, Multichannel Systems, Germany) were characterized using impedance spectroscopy. Impedance spectra in the frequency range 10 kHz to 1 MHz were obtained in phosphate buffer solution of physiological salinity at room temperature electrically connected by a silver-chloride ground electrode. A three-circuit model consisting of a resistor in series with capacitance ($C_{dl}$) and a resistor ($R_{t}$) in parallel was used to fit the spectra with $\chi^{2}\approx 0.2$. The resistance of the gap $R_{gap}$ was calculated from $R_{gap}=\rho_{s}/(5\pi d_{gap}$), with approximate values for the solution of conductivity $\rho_{s}=1$ $\Omega$m and $d_{gap}=1$ nm [104, 124]. The final values for the electrode model are $R_{\text{gap}}=6.37\times 10^{7}\,\Omega$, $C_{\text{dl}}=8.99\times 10^{-11}\,\text{F}$, and $R_{\text{t}}=7.92\times 10^{8}\,\Omega$.
@@ 6.4 Odor space analysis
We describe our method for selecting the most viable ORs for a given set of odorants in Algorithm 1. The method requires a response data matrix with entries (odor $\times$ OR), which we have retrieved from the DoOR 2.0.1 database [86, 24]; specifically, we use the provided door_response_matrix.csv, which contains responses of 78 sensory unit types to 692 odorants, normalized to [0,1][0,1][ 0 , 1 ] across all responding units. We first exclude all unit types not clearly linked to a single OR type, as described in [24], which leaves a total of 52 ORs to be analyzed. As only a subset of all unit type/odor combinations have been measured (about 14 %), we replace missing entries (NA) with zero response. This may introduce the issue that some odors are separated only by the absence of any response; in practice, this would make it impossible to distinguish such an odor from noise currents when no ligand is present. We mitigate this issue by including a ‘no ligand/SFR (spontaneous firing rate)’ category in the analysis, see below.
@@
Our algorithm for selecting odor/OR sets is easily implemented and performs the reduction quickly. Note however that the underlying dataset corresponds to response measurements of olfactory receptor neurons corrected for the baseline firing rate, and doesn’t cover cases where no ligand is present at all, as described above. Therefore, if the final selection given by our algorithm includes odors for which none of the selected ORs respond at all, this odor will be experimentally indistinguishable from no ligand binding. To cover this case, we have included a ‘no ligand’ category, i.e. a row in $D_{\text{all}}$ which corresponds to the spontaneous firing rate.
@@ 6.6 Design and Simulation of the Neuromorphic Hardware Platform
The system integrates an analog front-end with an event-driven, asynchronous spiking neural network and a microcontroller. It consists of: (i) a bioelectronic interface coupling lipid vesicles with synthetic olfactory receptors and microelectrode arrays, (ii) a low-noise analog amplifier for signal conditioning, (iii) a spiking neural network for energy-efficient sensor response processing, and (iv) a RISC-V microcontroller for spike decoding and synaptic weight configuration. Fig. 9 illustrates the layout of the full chip. The low-noise amplifier (LNA) exploits a folded cascode OTA (FC-OTA) with an active cascode and a flipped current follower (FCF) to improve signal integrity, transconductance efficiency, and noise performance. The FC-OTA topology enhances gain while maintaining a wide output swing, the active cascode increases output impedance for better gain stability, and the FCF reduces impedance at the folded node, ensuring efficient current transfer. A detailed overview of these techniques is provided, with further details in the work by R. Sanjay et al. [126]. Somatic and synaptic circuits are designed with weak-inversion transistors and follow the principles of well-known neuromorphic designs [108, 127] that have been demonstrated to scale to advanced technologies nodes [128, 129].
The schematic of the neuron circuit is illustrated in Fig. 7, block E. It uses a capacitor $C_{1}$ to represent the membrane potential $V_{mem}$. Three analog control voltages regulate the behavior of $V_{mem}$:
  • $V_{dc}$ sets the baseline charging current,
  • $V_{leak}$ defines the rate of decay, and
  • $V_{ref}$ determines the reset level during the refractory period.
This arrangement enables the hardware circuit to emulate a leaky integrate-and-fire (LIF) neuron. When $V_{mem}$ is below the firing threshold, its dynamics follow:
$V_{mem}(t+\Delta t)=\alpha V_{mem}(t)+I_{dc}\cdot\Delta t$ (2)
where $\alpha$ models the leakage over time. Unlike idealized models, here $\alpha$ depends on the membrane potential itself due to the circuit’s analog characteristics:
$\alpha=1-\frac{R_{leak}\cdot\Delta t}{V_{mem}}$ (3)
with $R_{leak}$ as the leakage rate (V/s), and $\Delta t$ as the simulation time step.
[…]
  1. 1.
    When a spike occurs, the neuron raises $!req$ to signal an event.
  2. 2.
    The receiver detects this and responds by asserting $!ack$.
  3. 3.
    Upon receiving acknowledgment, the neuron deasserts $!req$.
  4. 4.
    Finally, the receiver lowers $!ack$, completing the handshake.
This protocol ensures accurate and lossless communication of spikes, even in systems without a global clock.
Fig. 7, block D shows the synapse circuit. Synaptic inputs are implemented using a 4-bit digital representation. Each weight bit $W[i]$ (where $i=0\ldots 3$) defines the current strength, and its corresponding switch $S[i]$ enables or disables that contribution. When active, the current from each bit is summed to produce a total synaptic output current $i_{out}$:
$i_{out}=\sum_{i=0}^{3}S[i]\cdot W[i]$ (4)
This current is injected directly into the neuron’s $V_{mem}$ node, modifying its potential.
To balance the influence of bias (from $V_{dc}$) and synaptic input, transistor sizing in these branches is carefully chosen. The synaptic current is derived from subthreshold MOSFET operation, yielding an exponential current-voltage relationship. The discharge rate of the intermediate node $V_{sin}$ follows:
$R_{V_{sin}}=\frac{dV_{sin}}{dt}=\frac{I_{tot}}{C}=\frac{1}{C}\sum_{i}\frac{I_{ds,0_{i}}\cdot W_{i}}{L_{i}}e^{\frac{V_{W_{i}}-V_{th,n}}{nU_{T}}}$ (5)
Here, $V_{th,n}$ is the threshold voltage, $U_{T}=\frac{kT}{q}\approx 25$ mV at 300 K, and $n=1+\frac{C_{D}}{C_{ox}}$ is the subthreshold slope factor. The final synaptic current $i_{out}$ is produced via a P-type MOSFET $M_{0}$ and injected into the membrane node.